# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ Appendix G Modeling Assumptions and Limitations
This appendix outlines key assumptions in our modeling framework and discusses their implications and limitations.
G.1 User Traffic Modeling
Our model assumes that platform user traffic is driven entirely by human-created content, with AI-generated content affecting allocation but not the traffic volume. Formally, recall that the traffic function is given by $T(\mathbf{x})=\mu\cdot\left\|\mathbf{x}\right\|^{\gamma}$.
This assumption rests on two complementary observations. First, in many emerging or specialized domains, generative AI models are not yet capable of producing original content unless they are trained on relevant human data. Without such input, their ability to generalize meaningfully to new topics remains limited [34]. Second, users tend to value originality and authenticity, often showing a negative bias toward content they perceive as machine-generated [5, 13]. Together, these factors suggest that human contributions attract traffic, while AI-generated content plays a more supporting role.
Conceptually, we view human content as spanning the expressive boundaries of the content space. At the same time, GenAI fills the convex hull of this space with the ability to recreate, edit, and mix content it saw in its training set. We make the assumption that users are drawn to the creativity and novelty introduced by original human contributions (the "vertices" of the space), rather than its convex hull. We note that AI-generated content affects traffic distribution and utilities, as it competes with human content for attention.
Nonetheless, as generative technologies improve and their outputs become increasingly indistinguishable from human-created content [29, 25], this bias may diminish, and the presence of AI-generated content may begin to influence traffic more significantly.
G.2 Generative AI Quality Modeling
We model the quality of GenAI content as a function of the total shared creator data. Specifically, recall that we assume the following functional form on the AI-generated content quality: $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\cdot\sum_{i}x_{i}s_{i}$. This formulation reflects the intuition that in new or emerging topical domains, where the GenAI has not yet seen relevant data, each new high-quality data point can meaningfully improve the model’s ability to generate valuable content. Each such data point offers a fresh perspective, something that improves the GenAI’s understanding of the domain in a linear manner.
This assumption abstracts away several important effects. For instance, in mature domains, the marginal value of additional data may diminish, which is consistent with known scaling laws in machine learning [33]. We explore a generalization of this assumption in Appendix F, but most of our theoretical results rely on the linear case.
We also assume that GenAI content quality depends solely on human-shared data. In the early stages of learning about a novel topic, generative models cannot meaningfully bootstrap their understanding through self-generation. While recent advancements, such as chain-of-thought reasoning, allow models to generate internally consistent sequences, they still require initial human-grounded data to acquire domain knowledge [34]. Such capabilities may effectively increase $\alpha$, but they do not remove the fundamental dependence on human input.
G.3 Scalar Representation of Content Competitiveness
Each creator’s quality choice is represented as a single scalar $x_{i}\in[0,\infty)$, which embodies the overall “competitiveness” of their content. This quantity encompasses various dimensions of the creators’ produced body of content, including quality, quantity, trend alignment, expressiveness, domain specificity, and novelty.
@@ G.3 Scalar Representation of Content Competitiveness
This modeling choice is a common simplification in the literature of competitions on platforms, enabling meaningful analysis while preserving the essence of the competitive dynamics among creators [53, 52]. However, it’s worth noting that real-world scenarios may involve more intricate representations, such as continuous vectors of features (embeddings) [27]. Analyzing such representations requires entirely different tools, as it becomes less apparent which content is more competitive.
G.4 Continuous Data Sharing Decisions
@@
Our model enables creators to select their sharing level continuously, represented by $s_{i}\in[0,1]$. However, in practice, such granular control is rare, as platforms usually provide binary options, such as opting in or out.
We chose this design to facilitate flexibility and serve as a convenient starting point for future research that may generalize our findings to alternative modeling choices. Nevertheless, all our results remain valid under binary sharing choices. This is because Proposition 2 ensures that only binary sharing decisions are rational (excluding an edge case where all decisions are strategically identical, in which choosing one of the binary endpoints of the interval is an optimal choice).

### v2

@@ Appendix F Modeling Assumptions and Limitations
This section outlines key assumptions in our modeling framework and discusses their implications and limitations.
User Traffic Modeling
Our baseline model assumes that platform user traffic is driven entirely by human-created content, with AI-generated content affecting allocation but not the traffic volume. This encompasses the scenario of a fixed number of users as a special case when $\gamma=0$. In Section˜6, we relax this assumption by considering a traffic model that incorporates both human and AI-generated content, and show that our main results carry over with appropriately adjusted constants. More general traffic models–such as those capturing nonlinear interactions between human and AI quality–remain unexplored.
Generative AI Quality Modeling
We model the quality of GenAI content as a linear function of the total shared creator data: $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\cdot\sum_{i}x_{i}s_{i}$. In practice, the relationship between training data quantity and model quality typically exhibits diminishing returns, consistent with established scaling laws in machine learning [30]. To address this, Section˜6 considers a generalized quality function $Q_{\text{AI}}=\alpha(\sum_{i}x_{i}s_{i})^{\beta}$ with $\beta\in[0,1]$. While certain theoretical guarantees valid under the linear model do not hold in this generalized setting, our simulations suggest that the key qualitative insights persist. Other factors that may affect GenAI quality in practice–such as data diversity, representativeness, and complementarity among creators’ contributions–are not captured by our model and remain directions for future work.
Scalar Representation of Content Competitiveness
@@
Each creator’s quality choice is represented as a single scalar $x_{i}\in[0,\infty)$, embodying the overall “competitiveness” of their content. This quantity encompasses various dimensions, including quality, quantity, trend alignment, expressiveness, domain specificity, and novelty. This modeling choice is a common simplification in the literature [52, 50], enabling tractable analysis while preserving the essence of competitive dynamics. However, real-world scenarios may involve more intricate representations, such as continuous vectors of features [25], requiring entirely different analytical tools.
Continuous Data Sharing Decisions
@@
Our model enables creators to select their sharing level continuously, represented by $s_{i}\in[0,1]$. In practice, platforms usually provide binary options. Nevertheless, all our results remain valid under binary sharing choices, as Lemma˜2 ensures that only binary sharing decisions are rational.
Static Game Formulation
Our model captures a one-shot interaction between the platform and creators. In practice, creator-platform relationships are ongoing, and the quality of GenAI systems depends on data accumulated over time. Such dynamic settings introduce richer strategic considerations, including the timing of sharing decisions and the platform’s ability to update its allocation rule. Modeling these interactions as repeated or sequential games constitutes a natural extension of our framework.